40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Aug 25 15:43:29 2016, ericq, Update, SUS, OSEM issues SRM_OSEMS.pngPRM_LR.png
    Reply  Thu Aug 25 20:07:35 2016, gautam, Update, SUS, OSEM issues - maybe resolved? 
    Reply  Tue Sep 6 18:21:13 2016, rana, Update, SUS, OSEM issues PRMswing.png
       Reply  Tue Sep 6 20:52:42 2016, gautam, Update, SUS, OSEM issues P9060254.JPG
          Reply  Wed Sep 7 17:23:26 2016, gautam, Update, SUS, PRM LR fixed for now IMG_3208.JPGIMG_3211.JPG
             Reply  Tue Oct 11 00:43:50 2016, ericq, Update, SUS, PRM LR problematic again PRM_LR_Problem.pngsince_fix.pnglast10.png
                Reply  Tue Oct 11 10:15:04 2016, Steve, Update, SUS, PRM LR problematic again PRMalfoiled.png
                   Reply  Tue Oct 11 13:30:49 2016, gautam, Update, SUS, PRM LR problematic again PRM_LR_decay.png
                      Reply  Wed Oct 12 13:34:28 2016, gautam, Update, SUS, PRM LR problematic again 08.png
Message ID: 12475     Entry time: Tue Sep 6 20:52:42 2016     In reply to: 12472     Reply to this: 12476
Author: gautam 
Type: Update 
Category: SUS 
Subject: OSEM issues 

The modes look like they're at the right frequencies, so pointing more and more towards a LED or satellite box issue.

We peeked into the BS-PRM chamber via the ITMX chamber to see if we could shed any light on this situation. It's hard to get a picture that is in focus, but it looks quite clear that the LR LED (in the lower left when viewed from the HR side) isn't anywhere near as bright as the rest (see Attachment #1). Various hypothesis include failed LED / piece of Al foil blocking the LED / teflon aperture slipped over the LED. But looks like we can't solve this without opening up the BS-PRM chamber. The plan tomorrow is to open up the chamber, pull out the problematic coil. Once we have a better idea of what is going wrong, we can decide what the appropriate course of action is - replace the OSEM or something else. 


As part of the diagnosis, I switched the PRM and SRM satellite boxes earlier today evening around 6pm. They remain in this switched state for now.


Steve, we plan to take the BS-PRM heavy door off tomorrow morning.

Attachment 1: P9060254.JPG  1.302 MB  | Hide | Hide all
P9060254.JPG
ELOG V3.1.3-